how to make 1N soultion from 30% hydrogen peroxide?

Grade:12th pass

1 Answers

Vikas TU
6 years ago
Stock Solution is 30% means 30 g in 100ml 
first figure its ordinariness: 
Measure of Solute = (Normality X Volume X Equivalent Weight)/1000 
30g = (N X 100ml X 16 g/mol)/1000 
Typicality = 300/16 = 18.75 N 
N = 18.75 N 
Presently Stock arrangement : Required Solution 
N1V1 = N2V2 
(18.75N)V1 = (1N)(1000ml) 
V1 = 1000/18.75 = 53.33 mL 
Along these lines, take 53.33 mL of arrangement from stock i.e 30 % H2O2 and add water in it to make 1000 Ml. This will be 1N arrangement of H2O2.